OpenStudy (anonymous):

Solve: 3^2 - 2^2 / square root of 4 + square root of 49 using BIDMAS. Cheers

OpenStudy (mertsj):

\[\frac{3^3-2^2}{\sqrt{4}+\sqrt{49}}\]

OpenStudy (anonymous):

It is not written as a fraction

OpenStudy (anonymous):

You have to use BIDMAS in the process

OpenStudy (mertsj):

9-4/2+7= 9-2+7=7+7=14

OpenStudy (anonymous):

Why can't it be 9-9=0 if addition and subtraction according to BIDMAS is of equal priority?

OpenStudy (mertsj):

Because 4/2 means 4 divided by 2 and you must do the division before the addition and subtraction.

OpenStudy (anonymous):

what the monkey is a bidmas?

OpenStudy (mertsj):

brackets, indices, division multiplication, addition, subtraction

OpenStudy (anonymous):

oooh sort of like "please excuse my senile aunt sally"

OpenStudy (mertsj):

Precisely.

OpenStudy (anonymous):

That wasn't my question. I said why did you do subtraction before addition?

OpenStudy (anonymous):

If you did it the other way you would have got 9-9=0

OpenStudy (anonymous):

\[3^2 - 2^2 \div (\sqrt{4}) + \sqrt{49} \implies 9 - 4 \div 2 + 7\] According to BIDMAS rule, here you will divide first...

OpenStudy (anonymous):

I know that part but why subtract and then do the division when it's supposed to be the opposite. Can't it be 9-9=0?

OpenStudy (aravindg):

what is I in BIDMAS ?

OpenStudy (aravindg):

Oh its above :P ....i got the same doubt as satellite ..we usually use the term BODMAS

hartnn (hartnn):

it can't be 9-9 because the order of operation (evaluation) is from left to right. so, 9-2 will be evaluated first, not 2+7